OpenStudy (anonymous):

what is the limiting velocity of the object? m = 100kg g=9.8 m/sec^2 b=5kg/sec v(0)=10m/sec v(t) = ? m*(dv/dt)=mg-bv

OpenStudy (anonymous):

So you have a linear ODE: \[100\frac{dv}{dt}=100\cdot9.8-5v(t)\\ \frac{dv}{dt}+\frac{1}{20}v(t)=980\] Find the integrating factor: \[\large\mu(t)=e^{\int\frac{1}{20}~dt}=e^{\frac{1}{20}t}\] \[e^{1/20~t}\frac{dv}{dt}+\frac{1}{20}e^{1/20~t}v(t)=980e^{1/20~t}\] Note that the left side is a derivative: \[\frac{d}{dt}\bigg[e^{1/20~t}~v(t)\bigg]=980e^{1/20~t}\] Integrate both sides: \[e^{1/20~t}~v(t)=980\int e^{1/20~t}~dt\] And so on. Once you solve for \(v(t)\), the limiting velocity will be given by \(\displaystyle \lim_{t\to\infty}v(t)\).